Determine the relative extrema of the function h = x[sup]2[/sup] + (y - 3)[sup]2[/sup] + (z + 1)[sup]2[/sup]
1
Expert's answer
2011-10-18T11:32:40-0400
relative extrema: dh/dx=2x=0
=> x=0 is extremum point
dh/dy=2(y-3)=0
=> y=3 is extremum point
dh/dz=2(z+1)=0
=> z=-1 is extremum point
Hence (0,3,-1) is the relative extrema point of this function